OpenStudy (anonymous):

How to Solve? (x-2)^(3/4)=8

OpenStudy (anonymous):

\[(x-2)^{\frac{ 3 }{ 4 }}=8\] \[\sqrt[4]{(x-2)^3}=8\] \[(x-2)^3=8^4\] \[x-2=\sqrt[3]{4096}\] \[x=16+2\] \[x=18\]